Circular No. 5623

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET VÄISÄLÄ 1 (1992u) J. V. Scotti, Lunar and Planetary Laboratory, reports his recovery of

this comet with the Spacewatch telescope at Kitt Peak. The object is of stellar appearance and the position is in close agreement with the prediction on MPC 16379 (ephemeris on MPC 20653).

1992 UT  α_{2000}  δ_{2000}  m_{1}
Sept.26.47353  8^{h}00^{m}37.^{s}19  +14^{o}16′31′′.1
26.48174  8 00 37.80  +14 16 29.7  21.7
26.49755  8 00 39.31  +14 16 25.1
27.47478  8 02 07.37  +14 12 42.6
27.48235  8 02 07.98  +14 12 41.0  21.6
27.48815  8 02 08.56  +14 12 37.9

GRS 1734–292 E. Churazov, M. Gilfanov, B. Cordier, and M. C. Schmitz-Fraysse,

on behalf of the Granat team (Space Research Institute, Moscow; Service d’Astrophysique du CEA, Saclay; Centre d’Etude Spatiale des Rayonnements, Toulouse), report: “The Galactic Center region has been observed by Granat on Sept. 14, 15, 17, and 18. An outburst of hard x-ray emission from GRS 1734–292 (IAUC 5123) has been detected by the SIGMA telescope on Sept. 15 and 17. The 40- to 400-keV flux from the source rose from < 20 mCrab (2σ upper limit, Sept. 14) up to 36 mCrab on Sept. 15 and 17 (5.5σ detection), and then declined to < 20 mCrab on Sept. 18. The averaged Sept. 15 and 17 source spectrum was extremely hard, with a 40- to 400-keV power-law photon index of –0.6 ± 0.4. No evidence of a narrow annihilation line has been detected in the range 450-550 keV, with an upper limit on the narrow-line flux of 2 × 10^{−3} photon s^{−1} cm^{−2} (3σ).”

NOVA CYGNI 1992 D. Hanzl and E. Neureiterova report the following additional photom-

etry (cf. IAUC 5571, 5598) obtained at the N. Copernicus Observatory and Planetarium in Brno (comparison star HD 195664): Aug. 30.862 UT, V = 10.06 ± 0.07, B–V = –0.72 ± 0.08, U–B = –1.18 ± 0.10; Sept. 2.856, 9.91 ± 0.02, –0.64 ± 0.03, –1.17 ± 0.05; 17.834, 10.05 ± 0.06, –0.66 ± 0.05, –1.26 ± 0.03; 18.801, 10.04 ± 0.03, –0.64 ± 0.02, –1.25 ± 0.02.

1992 September 28 Daniel W. E. Green